# Does equivariance matter at scale?

### v1

@@ Abstract
Figure 1: Scaling with compute. The dots show the training compute budget and test loss in our experiments, the lines indicate the compute-optimal performance according to the scaling laws we find. The test losses of both non-equivariant ( ) and equivariant ( ) transformers scale as a power law with compute, and the equivariant model outperforms the non-equivariant model by a similar factor at all tested compute budgets.
Figure 2: Scaling with training data. We show the performance of the non-equivariant transformer ( ), non-equivariant transformer trained with data augmentation ( ), and equivariant transformer ( ) as a function of the number of unique tokens in the training dataset. All experiments use the same training compute budget. Equivariance improves data efficiency compared to the baseline, but data augmentation can close this gap.
@@ 1 Introduction
A common intuition is that strong inductive biases bring the biggest benefits when little training data is available, and that symmetry properties can just as well be learned from data given sufficient samples and compute. Recently, high-profile models of protein folding (Abramson et al., 2024) and conformer generation (Wang et al., 2023) have received considerable attention for their choice of non-equivariant architectures for geometric problems.
At the same time, there is reason to expect that equivariance is still beneficial in the large-data limit. Learning means successively narrowing down a hypothesis class based on evidence. From this perspective one can explain (Bahri et al., 2021) the empirical observation that test losses often scale as a power law with the training compute (Kaplan et al., 2020; Hoffmann et al., 2022). Whereas non-equivariant methods start from the space of virtually all functions, equivariant models start from the subspace of all functions that abide by the symmetries of the problem. The learning process may benefit from that by focusing solely on further refining this smaller hypothesis class, narrowing down to the correct solution with fewer training steps.
Until the theory of scaling laws is fully understood, the effects of equivariance on scaling is an empirical question, and in this work we study it empirically. We focus on a benchmark problem of modelling the physical interactions between rigid three-dimensional objects described by meshes. This task is known to be challenging (Allen et al., 2022). It is manifestly equivariant under $\mathrm{E}(3)$, the symmetry group of rotations, translations, and reflections. We compare a standard transformer architecture (Vaswani et al., 2017) to an $\mathrm{E}(3)$-equivariant transformer (Brehmer et al., 2023).
@@
Our experiments provide evidence for three conclusions. As expected, equivariance improves data efficiency. However, data augmentation largely closes this gap. Second, equivariant transformers are also more compute-efficient, and this advantage persists across all compute budgets studied. Both model classes exhibit power-law scaling behaviour. Finally, the optimal allocation of a training compute budget to model size and training steps differs between equivariant and non-equivariant models. Overall, our findings hint that strong inductive biases may not only yield benefits in the low-data regime, but can also be beneficial with large datasets and large compute budgets.
@@ 2 Background and related work
The scaling of neural network performance as a function of model size or training steps has been studied extensively (Ahmad & Tesauro, 1988; Hestness et al., 2017; Rosenfeld et al., 2019; Henighan et al., 2020). Kaplan et al. (2020) first observed that the test loss of autoregressive language models follows a power law over many orders of magnitude. Hoffmann et al. (2022) improved the methodology further and found the “Chinchilla” scaling laws, which still serve as a reference point for many language models. In our quantitative analysis of compute scaling, we largely follow their approach.
@@
There has been comparatively little research into the relation between inductive biases and scaling behaviour, perhaps because the transformer architecture (Vaswani et al., 2017) is so established in language modelling. Tay et al. (2022) compared the scaling behaviour of different architectures. Recently, Qiu et al. (2024) investigated how structured linear transformations in transformers affect scaling laws. The authors conclude that imposing structure in them can improve the scaling behaviour. Our work differs from both of these studies through its focus on symmetric problems and equivariant architectures.
@@
Equivariance has been found to improve performance, data efficiency, and robustness to out-of-domain generalization in fields as diverse as quantum mechanics and quantum field theory (Pfau et al., 2020; Hermann et al., 2020; Boyda et al., 2021; Gerdes et al., 2023), molecular force fields (Batatia et al., 2022; Batzner et al., 2022; Liao & Smidt, 2022; Musaelian et al., 2023; Batatia et al., 2023), generative models of molecules (Zeni et al., 2023; Igashov et al., 2024), particle physics (Bogatskiy et al., 2022; Gong et al., 2022; Spinner et al., 2024), biological and medical imaging (Veeling et al., 2018; Bekkers et al., 2018; Winkels & Cohen, 2018; Winkens et al., 2018; Mohamed et al., 2020; de Ruijter & Cesa, 2024; Suk et al., 2024), wireless communication (Hehn et al., 2024), and robotics (Wang et al., 2022a; b; c; Brehmer et al., 2024). The potential of equivariance to improve generalization has also been shown theoretically (Sokolic et al., 2017; Lyle et al., 2020; Elesedy & Zaidi, 2021; Sannai et al., 2021; Behboodi et al., 2022; Petrache & Trivedi, 2024).
@@ 3.1 Benchmark problem
Rigid-body modelling problem
We choose a rigid-body modelling problem as our benchmark. Three-dimensional meshes are initialized at some position, orientation, and velocity; they then interact with each other under gravity and collision forces. Concretely, the inputs to the network consist of a set of triangular meshes for two time points $t=t_{0},t_{0}+\Delta t$, and the task is to predict all mesh vertices at time $t=t_{0}+2\Delta t$. As a loss function and evaluation metric, we use the mean squared error of the predicted mesh vertex positions.
This problem satisfies all desiderata for our study. Rigid-body interactions are known to be challenging to model: collisions are difficult to detect, since they do not usually occur at or near vertices; the forces acting during a collision are nearly discontinuous (Bauza & Rodriguez, 2017; Pfrommer et al., 2021; Allen et al., 2022). Synthetic data can be generated cheaply with physics simulators. Finally, the physics of the process is clearly equivariant under $\mathrm{E}(3)$, provided that the direction of gravity is treated as a feature and rotated along with the scene.
@@ 3.2 Models
In selecting architectures, our main objective is not to achieve state-of-the-art results on the particular rigid-body benchmark problem we chose. That would lead us to highly problem-specific architectures (Allen et al., 2022; Rubanova et al., 2024). Instead, we aim for general-purpose architectures that are applicable to broad classes of problems.
@@
The transformer architecture (Vaswani et al., 2017) has become the de-facto standard across a wide range of machine learning tasks. It is versatile with respect to the input data, propagates gradients effectively, and scales well to large model sizes and input tokens. Most scaling studies have focused on transformers as well. We therefore use a standard pre-LN (Baevski & Auli, 2018) transformer with multi-query attention (Shazeer, 2019) as our non-equivariant architecture.
@@
Even this baseline architecture is hardly “free from inductive biases”. Because the tokens form not a sequence, but an unordered set, we do not use positional encoding. Therefore, the model is equivariant with respect to one of the symmetries of our problem: that of permutations of the input tokens. In this respect, there is no difference between the two architectures, and we do not compare to any models that are not permutation-equivariant.
@@
Second, we enforce object coherence and rigidity when computing the outputs. Either transformer model first outputs a translation vector and a rotation quaternion for each mesh face. These are averaged over each object, resulting in a translation vector and a rotation for each rigid object. These $\mathrm{E}(3)$ operations are then applied to the input meshes. In this way, the networks by design translate and rotate rigid objects consistently. We describe this procedure in more detail in Appendix B. In preliminary experiments, enforcing object rigidity in this way improved performance substantially compared to directly predicting the positions or velocities of mesh vertices. We also experimented with outputting and exponentiating elements of the Lie algebra for each object, but found that that worked marginally worse.
@@
Attention heads $2n$ $2n$
Scalars per key, query, value $64$ $8$
MV per key, query, value $2$
Hidden scalar channels in MLP $128n$ $8n$
@@
We tune the hyperparameters of both models manually. For both the baseline and equivariant transformer, we define a one-parameter family of hyperparameters, fixing the relation between the number of layers, attention heads, and channels to be linear. Our architectures are shown in Tbl. 1. Notably, we find that the equivariant transformer benefits from a more narrow architecture, which may be evidence of the expressivity of its multivector channels.
@@
We train all models with the Adam optimizer (Kingma, 2014), annealing the learning rate over the course of training from an initial value of $5\cdot 10^{-4}$ on a cosine schedule. For experiments with small FLOP budgets of less than $10^{18}$ nominal FLOPs, we find that this learning rate can be too small. This is in line with other works that find larger learning rates beneficial for smaller compute budgets (e. g.  Dubey et al., 2024). We therefore repeat these experiments with a higher learning rate of $10^{-3}$ or $2\cdot 10^{-3}$, depending on the compute budget, and report the better result. For simplicity, we use the same batch size of 64 samples (or on average $3.5\cdot 10^{5}$ tokens) for all experiments, even though this does not maximize GPU utilization and thus FLOP throughput. Early stopping is used in all experiments.
@@ 3.3 Scaling-law analysis
Second, we study the scaling with training data, fixing the training compute budget, the model size, and the number of training tokens. For both models we choose settings that performed compute-optimally in the first series of experiments for a compute budget of $10^{18}$ nominal FLOPs. The number of unique samples in the dataset is varied over five orders of magnitude, from $2\cdot 10^{6}$ tokens to $2\cdot 10^{11}$. The lower end of this scan corresponds to training for $6\cdot 10^{5}$ epochs, while every sample is seen only once on the upper end of this scan. For each of these settings, we train a baseline transformer, an equivariant transformer, and a baseline transformer trained with data augmentation, in which symmetry transformations are applied to the samples, independently for each epoch.
@@
The parameter $E$ represents the irreducible loss that even a perfect model cannot eliminate. Unlike in language or image modelling tasks, there is no clear reason to expect such an irreducible error of practically relevant size for the deterministic physics task we use as a benchmark. We treat the choice of whether to include $E$ as a fit parameter or fix it to zero as a hyperparameter and choose it through cross validation, as we will describe below.
@@
Scaling law Param. Baseline Equivariant
Central Lower Upper Central Lower Upper
Eq. (2): $\hat{L}(N,D)=A/{N^{\alpha}}+B/{D^{\beta}}$ $A$ $\mathbf{1.27}$ $0.484$ $5.07$ $\mathbf{0.000282}$ $0.000162$ $0.000607$
$B$ $\mathbf{0.202}$ $0.0108$ $0.361$ $\mathbf{469}$ $159$ $592$
$\alpha$ $\mathbf{0.909}$ $0.832$ $1.03$ $\mathbf{0.348}$ $0.293$ $0.417$
$\beta$ $\mathbf{0.379}$ $0.256$ $0.404$ $\mathbf{0.734}$ $0.689$ $0.747$
Eq. (4): $N^{*}(C)\propto C^{a}$ $a$ $\mathbf{0.294}$ $0.215$ $0.307$ $\mathbf{0.678}$ $0.619$ $0.711$
$b$ $\mathbf{0.706}$ $0.693$ $0.785$ $\mathbf{0.322}$ $0.289$ $0.381$
Eq. (5): $L^{*}(C)=F/{C^{\gamma}}$ $F$ $\mathbf{1.03}$ $0.124$ $1.89$ $\mathbf{0.14}$ $0.0524$ $0.517$
$\gamma$ $\mathbf{0.268}$ $0.213$ $0.284$ $\mathbf{0.236}$ $0.212$ $0.267$
@@ 4 Results
Figure 3: Test loss (dotted circles) and scaling-law predictions (background colour) as a function of model size and training tokens. Left: non-equivariant transformer. Right: equivariant transformer. In both cases, we observe good agreement of model performance and scaling-law fit.
@@ 4.1 Scaling with compute
and the exponents are compatible with each other within the confidence intervals shown in Tbl. 2. We visualize the empirical compute-loss measurements and the derived optimal compute-loss relationship in Fig. 2.
@@
Perhaps more surprising is that the optimal trade-off depends on the compute in a different way for the two models. We find that for a regular transformer, one should scale training tokens more steeply than model size. For the equivariant model, we find the opposite trend: one should put additional compute more in the model size than the training tokens. The compute-optimal model sizes thus become more similar for larger compute budgets.

### v2

@@ Abstract
Figure 1: Scaling with compute. The dots show the training compute budget and test loss in our experiments, the lines indicate the compute-optimal performance according to the scaling laws we find, the error bands estimate the uncertainty on the power-law coefficients. The test losses of both non-equivariant ( ) and equivariant ( ) transformers scale as a power law with compute, and the equivariant model outperforms the non-equivariant model by a similar factor at all tested compute budgets.
Figure 2: Scaling with training data. We show the performance of the non-equivariant transformer ( ), non-equivariant transformer trained with data augmentation ( ), and equivariant transformer ( ) as a function of the number of unique tokens in the training dataset. All experiments use the same training compute budget, which means that the number of epochs reduces from left to right. Equivariance improves data efficiency compared to the baseline, but data augmentation can close this gap.
@@ 1 Introduction
A common intuition is that strong inductive biases bring the biggest benefits when little training data is available, and that symmetry properties can just as well be learned from data given sufficient samples and compute. Recently, high-profile models of protein folding (Abramson et al., 2024) and conformer generation (Wang et al., 2023) have received considerable attention for their choice of non-equivariant architectures for geometric problems (Oxford Protein Informatics Group, 2024; Joshi, 2025; Twitter / X users, 2024).
At the same time, there is reason to expect that equivariance is still beneficial in the large-data limit. Learning means successively narrowing down a hypothesis class based on evidence. From this perspective one can explain (Bahri et al., 2024) the empirical observation that test losses often scale as a power law with the training compute (Kaplan et al., 2020; Hoffmann et al., 2022). Whereas non-equivariant methods start from the space of virtually all functions, equivariant models start from the subspace of all functions that abide by the symmetries of the problem. The learning process may benefit from that by focusing solely on further refining this smaller hypothesis class, narrowing down to the correct solution with fewer training steps.
Until the theory of scaling laws is fully understood, the effects of equivariance on scaling is an empirical question, and in this work we study it empirically. We focus on a benchmark problem of modeling the physical interactions between rigid three-dimensional objects described by meshes. This task is known to be challenging (Allen et al., 2023). It is manifestly equivariant under $\mathrm{E}(3)$, the symmetry group of rotations, translations, and reflections. We compare a standard transformer architecture (Vaswani et al., 2017) to an $\mathrm{E}(3)$-equivariant transformer (Brehmer et al., 2023).
@@
Our experiments provide evidence for three conclusions. As expected, equivariance improves data efficiency. However, data augmentation largely closes this gap. Second, equivariant transformers are also more compute-efficient, and this advantage persists across all compute budgets studied. Both model classes exhibit power-law scaling behavior. Finally, the optimal allocation of a training compute budget to model size and training steps differs between equivariant and non-equivariant models. Overall, our findings hint that strong inductive biases may not only yield benefits in the low-data regime, but can also be beneficial with large datasets and large compute budgets.
@@ 2 Background and related work
The scaling of neural network performance as a function of model size or training steps has been studied extensively (Ahmad & Tesauro, 1988; Hestness et al., 2017; Rosenfeld et al., 2020; Henighan et al., 2020). Kaplan et al. (2020) first observed that the test loss of autoregressive language models follows a power law over many orders of magnitude. Hoffmann et al. (2022) improved the methodology further and found the “Chinchilla” scaling laws, which still serve as a reference point for many language models. In our quantitative analysis of compute scaling, we largely follow their approach.
@@
There has been comparatively little research into the relation between inductive biases and scaling behavior, perhaps because the transformer architecture (Vaswani et al., 2017) is so established in language modeling. Tay et al. (2023) compared the scaling behavior of different architectures. Recently, Qiu et al. (2024) investigated how structured linear transformations in transformers affect scaling laws. The authors conclude that imposing structure in them can improve the scaling behavior. Our work differs from both of these studies through its focus on symmetric problems and equivariant architectures.
@@
Equivariance has been found to improve performance, data efficiency, and robustness to out-of-domain generalization in fields as diverse as quantum mechanics and quantum field theory (Pfau et al., 2020; Hermann et al., 2020; Boyda et al., 2021; Gerdes et al., 2023), molecular force fields (Batatia et al., 2022; Batzner et al., 2022; Liao & Smidt, 2023; Musaelian et al., 2023; Batatia et al., 2023), generative models of molecules (Zeni et al., 2025; Igashov et al., 2024), particle physics (Bogatskiy et al., 2022; Gong et al., 2022; Spinner et al., 2024), biological and medical imaging (Veeling et al., 2018; Bekkers et al., 2018; Winkels & Cohen, 2018; Winkens et al., 2018; Mohamed et al., 2020; de Ruijter & Cesa, 2024; Suk et al., 2024), wireless communication (Hehn et al., 2025), and robotics (Wang et al., 2022a; b; c; Brehmer et al., 2024). The potential of equivariance to improve generalization has also been shown theoretically (Sokolic et al., 2017; Lyle et al., 2020; Elesedy & Zaidi, 2021; Sannai et al., 2021; Behboodi et al., 2022; Petrache & Trivedi, 2024).
@@
On a side note, the second symmetry group that is relevant to the problem we study is that of permutations of the inputs. Both standard transformers (Vaswani et al., 2017) and GATr (Brehmer et al., 2023) are equivariant to it.
@@ 3.1 Benchmark problem
Rigid-body modeling problem
We choose a rigid-body modeling problem as our benchmark. Three-dimensional meshes are initialized at some position, orientation, and velocity; they then interact with each other under gravity and collision forces. Concretely, the inputs to the network consist of a set of triangular meshes for two time points $t_{0}$ and $t_{1}=t_{0}+\Delta t$, and the task is to predict all mesh vertices at time $t_{2}=t_{1}+\Delta t$. As a loss function and evaluation metric, we use the mean squared error of the predicted mesh vertex positions.
While this may sound like a problem straight from an undergraduate physics textbook, it serves as a microcosm of phenomena that appear throughout science and engineering: chaotic behavior in dynamical systems, extended 3D objects parameterized as meshes, and the symmetries of space and time. It also satisfies all desiderata for our study. Rigid-body interactions are known to be challenging to model: collisions are difficult to detect, since they do not usually occur at or near vertices; the forces acting during a collision are nearly discontinuous (Bauza & Rodriguez, 2017; Pfrommer et al., 2021; Allen et al., 2023). Synthetic data can be generated cheaply with physics simulators. Finally, the physics of the process is clearly equivariant under $\mathrm{E}(3)$, provided that the direction of gravity is treated as a feature and rotated along with the scene.
@@ 3.2 Models
In selecting architectures, our main objective is not to achieve state-of-the-art results on the particular rigid-body benchmark problem we chose. That would lead us to highly problem-specific architectures (Allen et al., 2023; Rubanova et al., 2024). Instead, we aim for general-purpose architectures that are applicable to broad classes of problems.
@@
The transformer architecture (Vaswani et al., 2017) has become the de-facto standard across a wide range of machine learning tasks. It is versatile with respect to the input data, propagates gradients effectively, and scales well to large model sizes and input tokens. Most scaling studies have focused on transformers as well. We therefore use a standard pre-LN (Baevski & Auli, 2019) transformer with multi-query attention (Shazeer, 2019) as our non-equivariant architecture.
@@
Even this baseline architecture is hardly “free from inductive biases.” Because the tokens form not a sequence, but an unordered set, we do not use positional encoding. Therefore, the model is equivariant with respect to one of the symmetries of our problem: that of permutations of the input tokens. In this respect, there is no difference between the two architectures, and we do not compare to any models that are not permutation-equivariant.
@@
As an aside, while we focus on the $\mathrm{E}(3)$ equivariance that the problem has when the direction of gravity is included as an input, one could alternatively treat the problem as $\mathrm{E}(2)$-symmetric in the plane orthogonal to the direction of gravity. We do not focus on that approach here, as the larger symmetry group $\mathrm{E}(3)$ is relevant in many problems and can easily be broken down to subgroups as needed.
@@
Second, we enforce object coherence and rigidity when computing the outputs. Either transformer model first outputs a translation vector and a rotation quaternion for each mesh face. These are averaged over each object, resulting in a translation vector and a rotation for each rigid object. These $\mathrm{E}(3)$ operations are then applied to the input meshes. In this way, the networks by design translate and rotate rigid objects consistently. We describe this procedure in more detail in Appendix B. In preliminary experiments, enforcing object rigidity in this way improved performance substantially compared to directly predicting the
@@
[…]
Scalars per key, query, value 646464 888
MV per key, query, value 222
[…]
@@
positions or velocities of mesh vertices. We also experimented with outputting and exponentiating elements of the Lie algebra for each object, but found that that worked marginally worse.
@@
We tune the hyperparameters of both models by manually varying the depth, width, and number of heads at low and intermediate compute budgets. For both the baseline and equivariant transformer, we define a one-parameter family of hyperparameters, fixing the relation between the number of layers, attention heads, and channels to be linear. Our architectures are shown in Tbl. 1. Notably, we find that the equivariant transformer benefits from a more narrow architecture, which may be evidence of the expressivity of its multivector channels.
@@
We train all models with the Adam optimizer (Kingma, 2015), annealing the learning rate over the course of training from an initial value of $5\cdot 10^{-4}$ on a cosine schedule. For experiments with small FLOP budgets of less than $10^{18}$ nominal FLOPs, we find that this learning rate can be too small. This is in line with other works that find larger learning rates beneficial for smaller compute budgets (e. g.  Dubey et al., 2024). We therefore repeat these experiments with a higher learning rate of $10^{-3}$ or $2\cdot 10^{-3}$, depending on the compute budget, and report the better result. For simplicity, we use the same batch size of 64 samples (or on average $3.5\cdot 10^{5}$ tokens) for all experiments, even though this does not maximize GPU utilization and thus FLOP throughput. Early stopping is used in all experiments.
@@ 3.3 Scaling-law analysis
Second, we study the scaling with training data, fixing the training compute budget, the model size, and the number of training tokens. For both models we choose settings that performed compute-optimally in the first series of experiments for a compute budget of $10^{18}$ nominal FLOPs. The number of unique samples in the dataset is varied over five orders of magnitude, from $2\cdot 10^{6}$ tokens to $2\cdot 10^{11}$. The lower end of this scan corresponds to training for $6\cdot 10^{5}$ epochs, while every sample is seen only once on the upper end of this scan. For each of these settings, we train three models: a baseline transformer, an equivariant transformer, and a baseline transformer trained with data augmentation. In the latter case, symmetry transformations are applied to the samples, independently for each epoch. Because the direction of gravity is not an explicit input to the transformer, the symmetry group consists of two-dimensional translations as well as rotations around the direction of gravity.
@@
Scaling law Param. Baseline Equivariant
Central Lower Upper Central Lower Upper
Eq. (2): $\hat{L}(N,D)=A/{N^{\alpha}}+B/{D^{\beta}}$ $A$ $\mathbf{1.27}$ 0.4840.4840.484 5.075.075.07 $\mathbf{0.000282}$ 0.0001620.0001620.000162 0.0006070.0006070.000607
$B$ $\mathbf{0.202}$ 0.01080.01080.0108 0.3610.3610.361 $\mathbf{469}$ 159159159 592592592
$\alpha$ $\mathbf{0.909}$ 0.8320.8320.832 1.031.031.03 $\mathbf{0.348}$ 0.2930.2930.293 0.4170.4170.417
$\beta$ $\mathbf{0.379}$ 0.2560.2560.256 0.4040.4040.404 $\mathbf{0.734}$ 0.6890.6890.689 0.7470.7470.747
Eq. (4): $N^{*}(C)\propto C^{a}$ $a$ $\mathbf{0.294}$ 0.2150.2150.215 0.3070.3070.307 $\mathbf{0.678}$ 0.6190.6190.619 0.7110.7110.711
$b$ $\mathbf{0.706}$ 0.6930.6930.693 0.7850.7850.785 $\mathbf{0.322}$ 0.2890.2890.289 0.3810.3810.381
Eq. (5): $L^{*}(C)=F/{C^{\gamma}}$ $F$ $\mathbf{1.03}$ 0.1240.1240.124 1.891.891.89 $\mathbf{0.14}$ 0.05240.05240.0524 0.5170.5170.517
$\gamma$ $\mathbf{0.268}$ 0.2130.2130.213 0.2840.2840.284 $\mathbf{0.236}$ 0.2120.2120.212 0.2670.2670.267
@@
The parameter $E$ represents the irreducible loss that even a perfect model cannot eliminate. Unlike in language or image modeling tasks, there is no clear reason to expect such an irreducible error of practically relevant size for the deterministic physics task we use as a benchmark. We treat the choice of whether to include $E$ as a fit parameter or fix it to zero as a hyperparameter and choose it through cross validation, as we will describe below.
@@ 4 Results
Figure 3: Test loss (dotted circles) and scaling-law predictions (background color) as a function of model size and training tokens. Left: non-equivariant transformer. Right: equivariant transformer. In both cases, we observe good agreement of model performance and scaling-law fit.
@@ 4.1 Scaling with compute
and the exponents are compatible with each other within the confidence intervals shown in Tbl. 2. We visualize the empirical compute–loss measurements and the derived optimal compute–loss relationship in Fig. 2.
@@
Perhaps more surprising is that the optimal trade-off depends on the compute in a different way for the two models. We find that for a regular transformer, one should scale training tokens more steeply than model
@@
size. For the equivariant model, we find the opposite trend: one should put additional compute more in the model size than the training tokens. The compute-optimal model sizes thus become more similar for larger compute budgets.